The perimeter of a rectangle is 68cm. The length is 4cm more than the width. Fine the dimention of the rectangle.
Answer by gems(12)

P = 68 cm
L = 4 + W
Formula:
P= 2L + 2W
68 = 2(4+w) + 2W
68 = 8 + 2W + 2W
68 - 8 = 4W
60 = 4W
15 = W
If W = 15
L = 4 + w
L = 4 + 15
L = 19
Therefore
the dimension is 15 cm by 19 cm